if A denotes the area bounded by f(x)=mod (sinx+cosx whole divided by x),x axis ,x=Π and x=3Π the A what r the max and min value of A

Grade:11

1 Answers

Jitender Singh IIT Delhi
10 years ago
Ans:
Using the inequality
m(b-a)\leq \int_{a}^{b}f(x)dx\leq M(b-a)
a = \pi , b = 3\pi
m\leq f(x)\leq M
Max. value of A
2\pi M
Min. value of A
2\pi m
You can easily find the m & M from f(x) (I am not finding b/c f(x) is not clear)
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty